OpenStudy (anonymous):

Let p1= (-2,7) and (5,-12). a. Find the slope between p1 and p2. b. Find an equation of the line that goes through the point p1 and p2. c. What is the x-intercept? d. What is the y-intercept?

OpenStudy (anonymous):

a. Find the slope between p1 and p2. slope = y2-y1/(x2-x1) b. Find an equation of the line that goes through the point p1 and p2. (y-y1) = slope* (x-x1) c. What is the x-intercept? Put y = 0 to find x intercept d. What is the y-intercept? y = mx + b ---- b is the y intercept